please help me figure this out
Solve for X 8(x-1)+3 = 7x-6-5x

Answer by stanbon(75887)

8(x-1)+3 = 7x-6-5x
8x-8+3=2x-6
8x-5=2x-6
6x=-1
x=-1/6
Cheers,
Stan H.